A parking meter takes only dimes and quarters. At the end of the day, the meter held $6.25 and contained a total of 46 coins. What was the number of dimes and quarters?
Found 2 solutions by richwmiller, TimothyLamb:
Answer by richwmiller(17219)

d+q=46,
25q+10d=625
d = 35, q = 11

Answer by TimothyLamb(4379)

10d + 25q = 625
d + q = 46
d = 46 - q
---
10d + 25q = 625
10(46 - q) + 25q = 625
460 - 10q + 25q = 625
15q = 165
Answer:
q = 11
d = 35
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php